OpenStudy (anonymous):

what is the smallest possible value of N which satisfy: ("==" means congruent) N==3(mod 4) N==2(mod 5) N==6(mod 7)

OpenStudy (mathmath333):

do u know chineese remainder theorem

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \huge \text{CRT}\hspace{.33em}\\~\\ N & \equiv 3\pmod 4\hspace{.33em}\\~\\ N & \equiv 2\pmod 5\hspace{.33em}\\~\\ N & \equiv 6\pmod 7\hspace{.33em}\\~\\ N& =\left[3\cdot (5\cdot7) \left((5\cdot7)^{-1}\pmod 4\right)\\~\\ +2\cdot (4\cdot7) \left((4\cdot7)^{-1}\pmod 5\right)\\~\\ +6\cdot (4\cdot5) \left((4\cdot5)^{-1}\pmod 7\right)\right]\pmod {4\cdot 5\cdot7}\\~\\ \end{align}}\)

OpenStudy (anonymous):

yes I do know CRT. I got 52 as my answer which is wrong. I actually didn't used the CRT, I took the two equations, N==2(mod 5) and N==6(mod7) writing them algebraically: N=5a+2=7b+6 which simplifies 5a=7b+6 taking modulo 5: 0==2b+1 now, subtracting 1 from both sides: 2b==0(mod5) b=5! 2*5=10==0(mod 5) plugging in the value of b in 7b+1: 7*5=35, +6 = 41, N==41(mod 35) AND the remaining one was: N==3(mod 4) applying the whole same method

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \huge \text{CRT}\hspace{.33em}\\~\\ N \equiv 3\pmod 4\hspace{.33em}\\~\\ N \equiv 2\pmod 5\hspace{.33em}\\~\\ N \equiv 6\pmod 7\hspace{.33em}\\~\\ N =\left[3\cdot (5\cdot7) \left((5\cdot7)^{-1}\pmod 4\right)\\~\\ +2\cdot (4\cdot7) \left((4\cdot7)^{-1}\pmod 5\right)\\~\\ +6\cdot (4\cdot5) \left((4\cdot5)^{-1}\pmod 7\right)\right]\pmod {4\cdot 5\cdot7}\\~\\ (5\cdot7)^{-1}\pmod 4\\~\\ 35x\equiv 1\pmod 4\\~\\ 3x\equiv 1\pmod 4\\~\\ \fbox{x=3}\\~\\ (4\cdot7)^{-1}\pmod 5\\~\\ 28y\equiv 1\pmod 5\\~\\ 3y\equiv 1\pmod 5\\~\\ \fbox{y=2}\\~\\ (4\cdot5)^{-1}\pmod 7\\~\\ 20z\equiv 1\pmod 7\\~\\ 6z\equiv 1\pmod 7\\~\\ \fbox{z=6}\\~\\ N =\left[3\cdot (5\cdot7) \left(3\right)\\~\\ +2\cdot (4\cdot7) \left(2\right)\\~\\ +6\cdot (4\cdot5) \left(6\right)\right]\pmod {4\cdot 5\cdot7}\\~\\ N =\dfrac{3\cdot (5\cdot7) (3)}{4\cdot 5\cdot7}\\~\\ +\dfrac{2\cdot (4\cdot7) (2)}{4\cdot 5\cdot7}\\~\\ +\dfrac{6\cdot (4\cdot5) (6)} {4\cdot 5\cdot7}\\~\\ =\dfrac{3\cdot (3)}{4}\\~\\ +\dfrac{2\cdot (2)}{5}\\~\\ +\dfrac{6\cdot) (6)} {7}\\~\\ =\dfrac{1}{4}\\~\\ +\dfrac{-1}{5}\\~\\ +\dfrac{1} {7}\\~\\ =\dfrac{35-28+20}{4\cdot5\cdot7}\\~\\ =\dfrac{27}{4\cdot5\cdot7}\\~\\ \Large =27 \end{align}}\)